# Soliton resolution for the energy-critical nonlinear heat equation in the radial case

### v2

@@ 1.1. Setting of the Problem
By [brezis-cazenave], given finite energy data, we let $T_{+}>0$ denote the maximal forward time of existence. Define the nonlinear energy functional associated with (1.1)
@@
and recall that by Caffarelli-Gidas-Spruck [CGS], all entire positive solutions to the stationary equation
@@ 1.2. Statement of the Main result
In the parabolic setting, similar results have been established for the harmonic map heat flow by Jendrej, Lawrie, and Schlag in [lawrie-harmonic-map, lawrie-harmonic-map-nonradial] following the bubbling theory for harmonic maps pioneered by Struwe [struwe], which was further developed in [topping, topping-annals, struwe, qing-tian, qing]. In fact, as explained in [kim2025classification], the connection between the harmonic map heat flow and the nonlinear heat equation can be seen by the following identity
@@
which relates the Laplacian operator for kk–equivariant functions on $\mathbb{R}^{2}$ to the one for radial functions on $\mathbb{R}^{2k+2}.$ Thus, we will often use the same language from the harmonic map heat flow literature. Finally, note that for the nonlinear heat equation, Theorem 1.1 has been conjectured to be true and, during the preparation of this paper, was also raised as an open question by Kim and Merle in [kim2025classification].
@@
The Soliton Resolution Conjecture states that the evolution of generic solutions to nonlinear dispersive equations decouples as a sum of modulated traveling waves (or Solitons) and a free radiation term. The conjecture arose in the 1970s following the numerical simulations of Fermi–Pasta–Ulam [pasta], Zabusky–Kruskal [zabusky]. Following the breakthrough work of Kenig and Merle [kenig-merle], the Soliton Resolution Conjecture has been established for some classes of dispersive PDEs. In particular, for the wave equation, the Soliton Resolution Conjecture has been proved in the radial case for either odd space dimensions or $D=4,6$ in [Duyckaerts-wave:d=3, Duyckaerts-wave-2, Duyckaerts-wave-5, Duyckaerts-wave-6, collot2022soliton] using the method of energy channels. On the other hand, using virial inequalities, Jia and Kenig in [jia-kenig] proved the sequential soliton resolution in dimension $D=6$. Recently, Jendrej and Lawrie in [lawrie-wave] established the Soliton Resolution Conjecture for the radial energy critical nonlinear wave equation in all space dimensions $D\geq 4$ using a novel argument based on the analysis of collision intervals. They also established the same result for the equivariant energy critical wave maps equation in [lawrie-wavemaps]. Finally, the Soliton Resolution Conjecture has also been established for the equivariant self-dual Chern–Simons–Schrödinger equation by Kim–Kwon–Oh in [oh-chern-simons].
@@
has been the subject of intensive study beginning with the seminal works of Giga and Kohn [giga-1, giga-2]. For a detailed introduction, see the excellent monograph of Quittner and Souplet [quitter]. Since our main theorem deals with the asymptotic behavior near the blow-up time, we briefly review some results in that direction. When the nonlinearity is energy subcritical, i.e., $1<p<\infty$ when $D=1,2$ and $1<p<\frac{D+2}{D-2}$ when $D\geq 3$, then [giga-1, giga-2] showed that any blow-up solution is of Type I, i.e., there exists a constant $C>0$ such that
@@
Otherwise, the blow-up is of Type II. In the energy critical setting, i.e., $p=\frac{D+2}{D-2},$ Filippas, Herrero and Velázquez [Filippas], established that the solution exhibits Type I blow-up if the initial data is positive and radially decreasing in dimension $D\geq 3.$ This result was later improved by Matano and Merle [matano2004nonexistence] where they proved, amongst other things, the same result with radial, positive and bounded initial data. Next, assuming that the initial datum is close to the ground state, Collot, Merle, and Raphaël in [collot-merle] established a Trichotomy in dimension $D\geq 7$; the solution either dissipates to zero, or approaches a rescaled Aubin–Talenti solution, or blows up in finite time, and the blow-up is of Type I. Thus, if the initial data is close to a ground state in dimension $D\geq 7$, the solution does not exhibit Type II blow-up. Recently, Wang and Wei in [wang2021refined] established that for positive initial datum in $L^{\infty}(\mathbb{R}^{D})$ and $D\geq 7$, any blow-up solution is of Type I. In contrast to the previous results, Theorem 1.1 is concerned with finite energy solutions that either exist globally in time or exhibit Type II blow-up. Examples of such solutions exhibiting non-trivial bubble tree decompositions have been recently constructed for the critical nonlinearity in dimensions $D\geq 7$ by del Pino, Musso, and Wei in [del_Pino_2021]. Furthermore, the recent work of Kim and Merle [kim2025classification] shows that under the assumption of radial symmetry, the above bubble tree constructions are the only possible examples in dimensions $D\geq 7.$
@@ 1.3. Summary of the Proof
Our proof is inspired by the recent breakthrough works of Jendrej–Lawrie, in particular, [lawrie-harmonic-map], where they established an analogous version of Theorem 1.1 for the harmonic map heat flow in the equivariant setting. However, we encounter several difficulties arising from the non-positivity of the nonlinear energy and slow decay of $\Lambda W$ [^1] Here $\Lambda=r\partial_{r}+\frac{D-2}{2}$ in lower dimensions $D<6$.
@@
On the other hand, the slow decay of $\Lambda W$ makes the modulation analysis more delicate in dimensions $D\in\{3,4,5\}$. The authors in [lawrie-wave] and [lawrie-wavemaps] circumvented this issue by performing a refined modulation analysis separately in dimensions $D\in\{4,5\}$. Instead of following their approach, we use our modified energy estimates, yielding a unified argument that works in all dimensions $D\geq 3$.
@@
The first step in the Jendrej–Lawrie framework is to establish a localized sequential compactness lemma, which essentially states that a sequence of functions with vanishing tension converges (locally in space) to a multi-bubble configuration. Thus, given a sequence of functions $u_{n}$ with bounded energy, a sequence $\rho_{n}\in(0,\infty)$ of scales, and tension $\mathcal{T}$ vanishing in $L^{2}$ relative to the scale $\rho_{n}$, i.e., $\lim_{n\to\infty}\rho_{n}\|\mathcal{T}(u_{n})\|_{L^{2}}=0$, there exists a subsequence of the $u_{n}$ that converges to a multi-bubble configuration up to large scales relative to $\rho_{n}$, i.e., $\lim_{n\to\infty}\boldsymbol{\delta}_{R_{n}\rho_{n}}(u_{n})=0$ for some sequence $R_{n}\to\infty$. Fortunately, sequential compactness for the nonlinear critical heat equation has been established recently in [lawrie-viet].
@@
The next step is to use Lemma 1.6 to prove the Soliton Resolution Conjecture along a sequence of times. This has been done using energy estimates in the context of the harmonic map heat flow in [lawrie-harmonic-map]. The key idea there was to establish localized energy monotonicity inequalities, which were used to deduce the size of the energy at later times in a suitable region. However, in our case, the localized nonlinear energy $\int_{0}^{\infty}\left(\frac{1}{2}|\nabla u|^{2}-\frac{1}{2^{*}}|u|^{2^{*}}\right)\phi^{2}r^{D-1}\,\mathrm{d}r$ is not necessarily non-negative, which prevents us from directly adapting the arguments in [lawrie-harmonic-map].
@@
We resolve this issue by using the radial Sobolev embedding, which gives us pointwise decay of the solution $u$ away from the origin. Thus localizing the modified energy $\tilde{E}(u)=\int_{0}^{\infty}\tilde{\mathbf{e}}(u)r^{D-1}\,\mathrm{d}r$ away from the origin, we deduce a localized energy monotonicity formula that is remarkably similar to the one derived for the harmonic map heat flow in [lawrie-harmonic-map]. However, this restriction on the localization region introduces more technical difficulties.
When $T_{+}=\infty$, we follow the arguments in [lawrie-harmonic-map] along with one new idea, which is required to show the positivity of the nonlinear energy away from the origin. To this end, we prove Lemma 2.10 where we show that the nonlinear energy of $u$ on the tail region $r_{0},\infty)$, for any $r_{0}>0$, is non-negative provided its $\mathcal{E}$–norm on $r_{0},\infty)$ is small enough. Thus, this lemma allows us to deduce the non-negativity of the nonlinear energy on tails, where we know that the solution $u$ has a small $\mathcal{E}$–norm since most of its energy is captured by the $N$-bubble configuration in some ball centered at the origin.
On the other hand, when $T_{+}<\infty$, the argument is more involved. This is because the energy of the solution $u$ is now given by the $N$-bubbles and a weak limit of the flow, denoted by $u^{*}$. Furthermore, we cannot localize the modified energy $\tilde{E}$ near the origin, and thus, we cannot access the bubbling region directly, forcing us to argue indirectly. This issue does not arise in the harmonic map heat flow case, where, for instance, one can directly show that the energy of the solution $u$ in the parabolic region is asymptotically equal to the sum of the energies of the $N$-bubbles. This fact, in turn, allows the authors in [lawrie-harmonic-map] to deduce that the energy of $u-u^{*}$ away from the parabolic region is small. Unfortunately, this line of reasoning does not work in our case, and thus, we are forced to argue in the reverse direction, which makes the argument more technical. See Section 3 for more details.
@@
Theorem 1.1 follows from the fact that $\lim_{t\to T_{+}}{\bf d}(t)=0$. To prove this, we argue by contradiction, i.e., suppose that ${\bf d}(t)$ does not converge to zero. Then, heuristically, this means that ${\bf d}(t)$ is large along a certain sequence of times, or, in other words, $u(t)$ deviates from the $N$-bubble configuration. Instead of analyzing this sequence of times, the key insight in [lawrie-wavemaps] is to consider a sequence of time intervals that keeps track of the dynamical history of the $N$-bubble configuration. Thus consider $a_{n},b_{n}]$, a sequence of time intervals where near the endpoints $a_{n}$ and $b_{n}$, $u(t)$ is close to an $N$-bubble configuration while inside $a_{n},b_{n}]$, $u(t)$ is close to an (NK)(N-K)-bubbles, where we choose $K$ to be the smallest non-negative integer such that the above properties hold. Observe that intuitively, $K\geq 1$ since otherwise, $u(t)$ will always be close to an $N$-bubble configuration, which would imply that $\lim_{t\to T_{+}}{\bf d}(t)=0$.
@@
Similar inequalities have been established in the context of nonlinear waves [lawrie-wave, lawrie-wavemaps] and the harmonic map heat flow [lawrie-harmonic-map]. However, the presence of a negative eigenvalue of the linearized operator (see Section 2.2) gives rise to an unstable direction, which makes our modulation analysis similar to that for the nonlinear wave equation. By controlling this unstable direction and adapting the arguments in [lawrie-wave], one can expect to deduce the inequality (1.19) when $D\geq 6$. We also expect that by using the refined modulation as in [lawrie-wave], one can push this down to $D\geq 4.$ However, it is not clear how to make this strategy work when $D=3.$ Thus, our last contribution is to prove Theorem 1.1 directly when $D\geq 3$ by getting around the technical difficulties arising from controlling the unstable direction in lower dimensions. To this end, we use energy estimates instead of differential inequalities for the unstable direction, thus avoiding any technical restrictions on the dimension, to prove (1.19). The key insight in this step is to realize that while the nonlinear heat equation has infinite speed of propagation, localized energy inequalities can be used to recover integral analogs of finite speed of propagation, similar to the nonlinear wave case. As a consequence, we can show that if a solution to (1.1) is close to an $N$-bubble configuration at some time in a region away from the origin (recall that our energy estimates work away from the origin) then this bound propagates for a short time implying that the distance to the $N$-bubble configuration still remains relatively small. This claim suffices to establish (1.19) when $D\geq 3$ without having to resort to estimates used in [lawrie-wave], where the authors proved a similar fact by controlling, among other things, the size of the unstable mode, which in turn required them to restrict the dimensions to $D\geq 6$ or resort to refined modulation analysis when $D\in\{4,5\}$.
@@ 2.1. Local Cauchy Theory
We first recall the local well-posedness theory for the heat equation in the energy space. See for instance, Theorem 1 in [brezis-cazenave], Theorem 2.1 in [roxanas], or Proposition 2.1 in [collot-merle, ikeda2025globaldynamicsenergycriticalnonlinear].
@@ 2.2. Multi-Bubble Configuration
Next, we recall some facts about solutions near multi-bubble configurations. For proofs and further references, see [lawrie-wave]. Denote $\mathcal{L}_{\mathcal{W}}$ the operator obtained by linearization of (1.1) about an $M$-bubble configuration $\mathcal{W}(\vec{\iota},\vec{\lambda})$
@@
We record here the following energy expansion around a multi-bubble configuration whose proof is similar to the proof of Lemma 2.15 in [lawrie-wave].
@@
Using a similar argument as in Lemma 2.17 in [lawrie-wave], when the value of the proximity function is small and the energy is close to the sum of energies of each bubble, we can find signs $\vec{\iota}$ and scales $\vec{\lambda}$ that realize the infimum in the above definition.
@@
Furthermore, similar to Lemma 2.20 in [lawrie-wave], if a function $w$ is close to two different multi-bubble configurations, then the scales of those two configurations are also the same up to a small constant.

### v3

@@ 1.1. Setting of the Problem
By [BC96], given finite energy data, we let $T_{+}>0$ denote the maximal forward time of existence. Define the nonlinear energy functional associated with (1.1)
@@
and recall that by Caffarelli-Gidas-Spruck [CGS89], all entire positive solutions to the stationary equation
@@ 1.2. Statement of the Main result
In the parabolic setting, similar results have been established for the harmonic map heat flow by Jendrej, Lawrie, and Schlag in [JL23a, JLS25] following the bubbling theory for harmonic maps pioneered by Struwe [STR85], which was further developed in [TOP97, TOP04, STR85, QT97, QIN95]. In fact, as explained in [KM25], the connection between the harmonic map heat flow and the nonlinear heat equation can be seen by the following identity
@@
which relates the Laplacian operator for kk–equivariant functions on $\mathbb{R}^{2}$ to the one for radial functions on $\mathbb{R}^{2k+2}.$ Thus, we will often use the same language from the harmonic map heat flow literature. Finally, note that for the nonlinear heat equation, Theorem 1.1 has been conjectured to be true and, during the preparation of this paper, was also raised as an open question by Kim and Merle in [KM25].
@@
The Soliton Resolution Conjecture states that the evolution of generic solutions to nonlinear dispersive equations decouples as a sum of modulated traveling waves (or Solitons) and a free radiation term. The conjecture arose in the 1970s following the numerical simulations of Fermi–Pasta–Ulam [FPU55], Zabusky–Kruskal [ZK65]. Following the breakthrough work of Kenig and Merle [KM06], the Soliton Resolution Conjecture has been established for some classes of dispersive PDEs. In particular, for the wave equation, the Soliton Resolution Conjecture has been proved in the radial case for either odd space dimensions or $D=4,6$ in [DKM12, DKM13, DKM23, DKM+22, CDK+24] using the method of energy channels. On the other hand, using virial inequalities, Jia and Kenig in [JK17] proved the sequential soliton resolution in dimension $D=6$. Recently, Jendrej and Lawrie in [JL23b] established the Soliton Resolution Conjecture for the radial energy critical nonlinear wave equation in all space dimensions $D\geq 4$ using a novel argument based on the analysis of collision intervals. They also established the same result for the equivariant energy critical wave maps equation in [JL22]. Finally, the Soliton Resolution Conjecture has also been established for the equivariant self-dual Chern–Simons–Schrödinger equation by Kim–Kwon–Oh in [KKO22].
@@
has been the subject of intensive study beginning with the seminal works of Giga and Kohn [GK87, GMS04]. For a detailed introduction, see the excellent monograph of Quittner and Souplet [QS19]. Since our main theorem deals with the asymptotic behavior near the blow-up time, we briefly review some results in that direction. When the nonlinearity is energy subcritical, i.e., $1<p<\infty$ when $D=1,2$ and $1<p<\frac{D+2}{D-2}$ when $D\geq 3$, then [GK87, GMS04] showed that any blow-up solution is of Type I, i.e., there exists a constant $C>0$ such that
@@
Otherwise, the blow-up is of Type II. In the energy critical setting, i.e., $p=\frac{D+2}{D-2},$ Filippas, Herrero and Velázquez [FHV00], established that the solution exhibits Type I blow-up if the initial data is positive and radially decreasing in dimension $D\geq 3.$ This result was later improved by Matano and Merle [MM04] where they proved, amongst other things, the same result with radial, positive and bounded initial data. Next, assuming that the initial datum is close to the ground state, Collot, Merle, and Raphaël in [CMR17] established a Trichotomy in dimension $D\geq 7$; the solution either dissipates to zero, or approaches a rescaled Aubin–Talenti solution, or blows up in finite time, and the blow-up is of Type I. Thus, if the initial data is close to a ground state in dimension $D\geq 7$, the solution does not exhibit Type II blow-up. Recently, Wang and Wei in [WW21] established that for positive initial datum in $L^{\infty}(\mathbb{R}^{D})$ and $D\geq 7$, any blow-up solution is of Type I. In contrast to the previous results, Theorem 1.1 is concerned with finite energy solutions that either exist globally in time or exhibit Type II blow-up. Examples of such solutions exhibiting non-trivial bubble tree decompositions have been recently constructed for the critical nonlinearity in dimensions $D\geq 7$ by del Pino, Musso, and Wei in [dMW21]. Furthermore, the recent work of Kim and Merle [KM25] shows that under the assumption of radial symmetry, the above bubble tree constructions are the only possible examples in dimensions $D\geq 7.$
@@ 1.3. Summary of the Proof
Our proof is inspired by the recent breakthrough works of Jendrej–Lawrie, in particular, [JL23a], where they established an analogous version of Theorem 1.1 for the harmonic map heat flow in the equivariant setting. However, we encounter several difficulties arising from the non-positivity of the nonlinear energy and slow decay of $\Lambda W$ [^1] Here $\Lambda=r\partial_{r}+\frac{D-2}{2}$ in lower dimensions $D<6$.
@@
On the other hand, the slow decay of $\Lambda W$ makes the modulation analysis more delicate in dimensions $D\in\{3,4,5\}$. The authors in [JL23b] and [JL22] circumvented this issue by performing a refined modulation analysis separately in dimensions $D\in\{4,5\}$. Instead of following their approach, we use our modified energy estimates, yielding a unified argument that works in all dimensions $D\geq 3$.
@@
The first step in the Jendrej–Lawrie framework is to establish a localized sequential compactness lemma, which essentially states that a sequence of functions with vanishing tension converges (locally in space) to a multi-bubble configuration. Thus, given a sequence of functions $u_{n}$ with bounded energy, a sequence $\rho_{n}\in(0,\infty)$ of scales, and tension $\mathcal{T}$ vanishing in $L^{2}$ relative to the scale $\rho_{n}$, i.e., $\lim_{n\to\infty}\rho_{n}\|\mathcal{T}(u_{n})\|_{L^{2}}=0$, there exists a subsequence of the $u_{n}$ that converges to a multi-bubble configuration up to large scales relative to $\rho_{n}$, i.e., $\lim_{n\to\infty}\boldsymbol{\delta}_{R_{n}\rho_{n}}(u_{n})=0$ for some sequence $R_{n}\to\infty$. Fortunately, sequential compactness for the nonlinear critical heat equation has been established recently in [LAW23].
@@
The next step is to use Lemma 1.6 to prove the Soliton Resolution Conjecture along a sequence of times. This has been done using energy estimates in the context of the harmonic map heat flow in [JL23a]. The key idea there was to establish localized energy monotonicity inequalities, which were used to deduce the size of the energy at later times in a suitable region. However, in our case, the localized nonlinear energy $\int_{0}^{\infty}\left(\frac{1}{2}|\nabla u|^{2}-\frac{1}{2^{*}}|u|^{2^{*}}\right)\phi^{2}r^{D-1}\,\mathrm{d}r$ is not necessarily non-negative, which prevents us from directly adapting the arguments in [JL23a].
@@
We resolve this issue by using the radial Sobolev embedding, which gives us pointwise decay of the solution $u$ away from the origin. Thus localizing the modified energy $\tilde{E}(u)=\int_{0}^{\infty}\tilde{\mathbf{e}}(u)r^{D-1}\,\mathrm{d}r$ away from the origin, we deduce a localized energy monotonicity formula that is remarkably similar to the one derived for the harmonic map heat flow in [JL23a]. However, this restriction on the localization region introduces more technical difficulties.
When $T_{+}=\infty$, we follow the arguments in [JL23a] along with one new idea, which is required to show the positivity of the nonlinear energy away from the origin. To this end, we prove Lemma 2.10 where we show that the nonlinear energy of $u$ on the tail region $r_{0},\infty)$, for any $r_{0}>0$, is non-negative provided its $\mathcal{E}$–norm on $r_{0},\infty)$ is small enough. Thus, this lemma allows us to deduce the non-negativity of the nonlinear energy on tails, where we know that the solution $u$ has a small $\mathcal{E}$–norm since most of its energy is captured by the $N$-bubble configuration in some ball centered at the origin.
On the other hand, when $T_{+}<\infty$, the argument is more involved. This is because the energy of the solution $u$ is now given by the $N$-bubbles and a weak limit of the flow, denoted by $u^{*}$. Furthermore, we cannot localize the modified energy $\tilde{E}$ near the origin, and thus, we cannot access the bubbling region directly, forcing us to argue indirectly. This issue does not arise in the harmonic map heat flow case, where, for instance, one can directly show that the energy of the solution $u$ in the parabolic region is asymptotically equal to the sum of the energies of the $N$-bubbles. This fact, in turn, allows the authors in [JL23a] to deduce that the energy of $u-u^{*}$ away from the parabolic region is small. Unfortunately, this line of reasoning does not work in our case, and thus, we are forced to argue in the reverse direction, which makes the argument more technical. See Section 3 for more details.
@@
Theorem 1.1 follows from the fact that $\lim_{t\to T_{+}}{\bf d}(t)=0$. To prove this, we argue by contradiction, i.e., suppose that ${\bf d}(t)$ does not converge to zero. Then, heuristically, this means that ${\bf d}(t)$ is large along a certain sequence of times, or, in other words, $u(t)$ deviates from the $N$-bubble configuration. Instead of analyzing this sequence of times, the key insight in [JL22] is to consider a sequence of time intervals that keeps track of the dynamical history of the $N$-bubble configuration. Thus consider $a_{n},b_{n}]$, a sequence of time intervals where near the endpoints $a_{n}$ and $b_{n}$, $u(t)$ is close to an $N$-bubble configuration while inside $a_{n},b_{n}]$, $u(t)$ is close to an (NK)(N-K)-bubbles, where we choose $K$ to be the smallest non-negative integer such that the above properties hold. Observe that intuitively, $K\geq 1$ since otherwise, $u(t)$ will always be close to an $N$-bubble configuration, which would imply that $\lim_{t\to T_{+}}{\bf d}(t)=0$.
@@
Similar inequalities have been established in the context of nonlinear waves [JL23b, JL22] and the harmonic map heat flow [JL23a]. However, the presence of a negative eigenvalue of the linearized operator (see Section 2.2) gives rise to an unstable direction, which makes our modulation analysis similar to that for the nonlinear wave equation. By controlling this unstable direction and adapting the arguments in [JL23b], one can expect to deduce the inequality (1.19) when $D\geq 6$. We also expect that by using the refined modulation as in [JL23b], one can push this down to $D\geq 4.$ However, it is not clear how to make this strategy work when $D=3.$ Thus, our last contribution is to prove Theorem 1.1 directly when $D\geq 3$ by getting around the technical difficulties arising from controlling the unstable direction in lower dimensions. To this end, we use energy estimates instead of differential inequalities for the unstable direction, thus avoiding any technical restrictions on the dimension, to prove (1.19). The key insight in this step is to realize that while the nonlinear heat equation has infinite speed of propagation, localized energy inequalities can be used to recover integral analogs of finite speed of propagation, similar to the nonlinear wave case. As a consequence, we can show that if a solution to (1.1) is close to an $N$-bubble configuration at some time in a region away from the origin (recall that our energy estimates work away from the origin) then this bound propagates for a short time implying that the distance to the $N$-bubble configuration still remains relatively small. This claim suffices to establish (1.19) when $D\geq 3$ without having to resort to estimates used in [JL23b], where the authors proved a similar fact by controlling, among other things, the size of the unstable mode, which in turn required them to restrict the dimensions to $D\geq 6$ or resort to refined modulation analysis when $D\in\{4,5\}$.
@@ 2.1. Local Cauchy Theory
We first recall the local well-posedness theory for the heat equation in the energy space. See for instance, Theorem 1 in [BC96], Theorem 2.1 in [GR18], or Proposition 2.1 in [CMR17, INP24].
@@ 2.2. Multi-Bubble Configuration
Next, we recall some facts about solutions near multi-bubble configurations. For proofs and further references, see [JL23b]. Denote $\mathcal{L}_{\mathcal{W}}$ the operator obtained by linearization of (1.1) about an $M$-bubble configuration $\mathcal{W}(\vec{\iota},\vec{\lambda})$
@@
We record here the following energy expansion around a multi-bubble configuration whose proof is similar to the proof of Lemma 2.15 in [JL23b].
@@
Using a similar argument as in Lemma 2.17 in [JL23b], when the value of the proximity function is small and the energy is close to the sum of energies of each bubble, we can find signs $\vec{\iota}$ and scales $\vec{\lambda}$ that realize the infimum in the above definition.
@@
Furthermore, similar to Lemma 2.20 in [JL23b], if a function $w$ is close to two different multi-bubble configurations, then the scales of those two configurations are also the same up to a small constant.